OpenStudy (anonymous):

Question 8 (Multiple Choice Worth 1 points) If f(x) = x-3/3 , then what is f–1(x) ?

OpenStudy (anonymous):

We write like this. y=(x-3)/3 and we need to get x rather than y. So y=x/3 -1 => x/3=y+1 => x=3y+3 So f(y)=3y+3. Or f-1(x)=3x+3. Hope it helped.

OpenStudy (anonymous):

very much thankyou (::

OpenStudy (anonymous):

Welcome